Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 8:24 Friday 16 July 2021 (17549) Print this report
Oplev hardware

With Hirata-san. 

Yesterday we installed the SR2 oplev hardware. We routed a power cable from the rack to the SLED. 

After turning on the SLED we confirmed that the light reaches the upper set of optics not far away from a suitable position. 

The next step is to check that the beam is reflected by the suspended mirror close to its centre. 

Comments to this report:
fabian.arellano - 17:45 Monday 19 July 2021 (17583) Print this report

With Hirata-san.

  • Using the horizontal fishng rods I moved the IP close to the setpoint.
  • The GAS filters were already close to the setpoint.
  • I open the vacuum chamber and aligned the collimator such that the beam reached the suspended mirror close to the centre.
  • The I aligned the receiving optics such  that the beams reached both QPDs.
  • I made sure the beam in the length sensing path went through the lens at the centre.

Tomorrow I will check the position of the micrometers and aim to set the lenth sensing QPD in a suitable place with respect to the lens.

Hirata-san checked the new screw system for locking the suspended breadboard.

fabian.arellano - 16:18 Tuesday 20 July 2021 (17592) Print this report
  • I moved the vertical micrometer stages to around 6 mm, which is close to their midrange and adjusted the two folding mirrors for the beams reach the QPD centers.
  • I set the length sensing QPD at the appropriate position with respect to the lens:
    • I applied white noise actuation to TM-P and observed the amplitude of the 0.843 Hz resonance peak in the amplitude spectral density of the LEN VER  channel.
    • I normalized the channel values at the resonance with TILT VER in order to cancel out variations in TM-P actuation at each ASD measurement.
    • The plot shows the values before and after normalization.

As can be seen in the plot, the position of the length sensing QPD that minimizes the coupling from TM-P is at 8.50 mm. The ASD plot shows that the LEN VER (in blue) is at the noise floor level at such position.

Images attached to this comment
Non-image files attached to this comment
fabian.arellano - 16:27 Wednesday 21 July 2021 (17614) Print this report

Calibration of the oplev QPD readout in terms of the beam displacement

Today I used Miyo-kun's script to acquire the data. The current location of the data is 


The files are


I'll analyze the data and post the results later.

fabian.arellano - 15:36 Thursday 05 August 2021 (17816) Print this report

I analyzed  the calibration data.

The results are shown in the following table:

QPD parameter calibration factor (QPD-du/mm) Linear range (mm) Tota range (µm)
LEN VER -8.64 [ 5.40  , 5.54 ] ±70
LEN HOR -8.02 [ 8.48 ,  8.63 ] ±75
TILT VER -1.53 [ 5.6  , 6.35 ] ±375
TILT HOR -1.31 [ 9.05 , 10.0 ] ±475


  • The linear ranges for the Tilt QPD are larger than those for the length QPD.
  • The readout of the QPDs is not in units of counts. Counts refers to the DAC and ADC. I call the QPD readout units QPD displacement units or QPD-du.
  • I fitted the data in Excel and not with the script provided in the medm screen because I wanted to choose the data points considered to be within the linear range and report the size of the range.
Images attached to this comment
Non-image files attached to this comment
terrence.tsang - 14:27 Wednesday 24 November 2021 (18980) Print this report

Using Fabian's data measured in 17614, I've obtained the calibration factors as follows.
Calibration notebooks can be found in /kagra/Dropbox/Subsystems/VIS/vis_commissioning/sr2/notebook/calibration/ 

calibration factor: -115.477 µm/d.u.
linear range (5% nonlinearity):  ±67.785 µm

calibration factor: -106.598 µm/d.u.
linear range (5% nonlinearity): ±62.733 µm

calibration factor: -706.673 µm/d.u.
linear range (5% nonlinearity): ±407.397 µm

calibration factor: -606.410 µm/d.u.
linear range (5% nonlinearity): ±338.377 µm

These calibration factors are implemented in K1:VIS-SR2_TM_OPLEV_{TILT, LEN}_{VER, HOR}_GAIN.

Images attached to this comment
Search Help